OpenStudy (anonymous):

Which property of real numbers is illustrated by the following statement? a + m k = a + k m

OpenStudy (anonymous):

A+B = B+A commutative property over addition.. A*B=B*A commutative property over multiplication

OpenStudy (anonymous):

so it is commutative property of multiplication

OpenStudy (anonymous):

yes